# Provable Sample-Efficient Transfer Learning Conditional Diffusion Models via Representation Learning

### v1

@@ A.4 Auxiliary Lemmas
$\displaystyle\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big{|}\lesssim\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK}}+\frac{b\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK},$ (A.120)
$\displaystyle\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\varphi(x_{i}^{k}))^{2}-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi^{2}]\Big{|}\lesssim b\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK}}+\frac{b^{2}\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK}.$ (A.121)
@@
$\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big{|}\lesssim\mathcal{R}_{K,n}(\Phi^{(j_{0}:)})+\sqrt{\frac{(B\epsilon_{j_{0}}+B_{0})\log\big{(}(\log nK)/\delta\big{)}}{nK}}+\frac{b\log\big{(}(\log nK)/\delta\big{)}}{nK}.$ (A.122)
From now on we reason on the conjunction of (A.120), (A.121) and (A.122). Define
$U_{j}=B\epsilon_{j}+B_{0}+b\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK}}+\frac{b^{2}\log\big{(}\log(b/\epsilon_{j})/\delta\big{)}}{nK}.$ (A.123)
and thus for any $\varphi\in\Phi^{(j)}$, we have $\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\varphi(x_{i}^{k}))^{2}\leq CU_{j}$ for some absolute constant $C$ by (A.121), indicating that $\mathcal{R}_{K,n}(\Phi^{(j)})\leq\phi_{K,n}(CU_{j})\leq\sqrt{C}\phi_{K,n}(U_{j})$. For any $j\leq j_{0}$,
$U_{j}\leq 2(B\epsilon_{j}+B_{0})+b\sqrt{C}\phi_{K,n}(U_{j})+\frac{2b^{2}\log\big{(}(\log nK)/\delta\big{)}}{nK}.$ (A.124)
@@
$U_{j}\lesssim b^{2}r_{K,n}^{*}+B\epsilon_{j}+B_{0}+\frac{b^{2}\log\big{(}(\log nK)/\delta\big{)}}{nK}=:r_{K,n}(\epsilon_{j}).$ (A.125)
Therefore, for any $\varphi\in\Phi^{(j)},j\leq j_{0}$, by (A.120),
$\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big{|}$ $\lesssim\phi_{K,n}(r_{K,n}(\epsilon_{j}))+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big{(}(\log nK)/\delta\big{)}}{nK}}+\frac{b\log\big{(}(\log nK)/\delta\big{)}}{nK}$ (A.126)
@@
$\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big{|}\lesssim F_{K,n}(\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]).$ (A.127)
@@
$\displaystyle\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big{(}(\log nK)/\delta\big{)}}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log nK)/\delta\big{)}}{nK}},$ (A.128)
$\displaystyle\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})\leq 2\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big{(}(\log nK)/\delta\big{)}}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big{(}(\log nK)/\delta\big{)}}{nK}}.$ (A.129)
Here $C^{\prime}$ is an absolute constant. Moreover, when $\varphi\in\Phi^{(j)}$ for $j>j_{0}$, we have $\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{b}{nK}$, and according to (A.122),
$\Big{|}\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big{|}\lesssim F_{K,n}(\varepsilon_{j_{0}}).$ (A.130)
@@
A.5 Verifying Task Diversity Assumption
@@
Proposition A.13.
Suppose $\mathcal{F}=\textbf{conv}(\mathcal{F})$, and $\inf_{x,y}\frac{p^{k}(x,y)}{p^{0}(x,y)}\geq\lambda_{k}$ for any $1\leq k\leq K$. Let $\lambda=\sum_{k=1}^{K}\lambda_{k}$. Then $\mathbb{P}^{1},\cdots,\mathbb{P}^{K}$ are $(\widetilde{\nu},\widetilde{\Delta})$-diverse over $\mathbb{P}^{0}$ with $\widetilde{\nu}=\lambda/(2K)$,
$\widetilde{\Delta}=2\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}\mathbb{E}_{t,x_{t}}\left[\left\|\frac{1}{\lambda}\sum_{k=1}^{K}\lambda_{k}\nabla\log p_{t}^{k}(x_{t}|y)-\nabla\log p_{t}^{0}(x_{t}|y)\right\|^{2}\right].$ (A.131)
@@
$L^{\mathbb{P}^{0}}(s_{\widetilde{f}^{0},h})$ $=\mathbb{E}_{\mathbb{P}^{0}}\left[\|\widetilde{f}^{0}(x_{t},h(y),t)-\nabla\log p_{t}^{0}(x_{t}|y)\|^{2}\right]$ (A.132)
@@
Appendix B Proofs in Section 4
B.1 Proofs of Score Network Approximation
Theorem B.1 (Thm. 4.1).
@@ B.1 Proofs of Score Network Approximation
[…]
@@
$L_{f}=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),W_{f}=\mathcal{O}\left(\frac{\log^{3(d_{x}+d_{y})/2}(nK/(\varepsilon\delta))}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right),$ (B.3)
@@
$L_{h}=\mathcal{O}\left(\log(1/\varepsilon)\right),W_{h}=\mathcal{O}\left(\varepsilon^{-D_{y}}\log(1/\varepsilon)\right),$ (B.4)
@@
$\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ (B.5)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}]$ $\leq\mathbb{E}_{t,x_{t},y}\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}\cdot\mathbbm{1}_{\|x_{t}\|\leq R_{f}}$ (B.6)
@@
where $\Omega_{R_{f}}=[-R_{f},R_{f}]^{d_{x}}\times[0,1]^{d_{y}}\times[T_{0},T]$. By Proposition B.2, B.3,
$\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ (B.7)
$\leq\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\mathbb{E}_{y\sim\mathbb{P}}[2\gamma_{f}^{2}\|h(y)-h_{*}(y)\|^{2}]+16M_{f}^{2}\exp(-C_{1}^{\prime}R_{f}^{2})$
@@
$\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h_{*}(y),t)-f^{\mathbb{P}}(x_{t},h_{*}(y),t)\|^{2}]\leq d_{x}\varepsilon^{2},$ (B.8)
@@
$L_{f}=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),W_{f}=\mathcal{O}\left(\frac{\log^{3(d_{x}+d_{y})/2}(nK/(\varepsilon\delta))}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right),$ (B.9)
@@
By Lemma A.2 and B.6, $f_{*}$ is $\gamma_{1}$-Lipschitz in $x$, $\gamma_{2}$-Lipschitz in $w$, and $\gamma_{3}$-Lipshcitz in $t$ in a bounded domain $\Omega_{R_{f}}$, where $\gamma_{1}=C_{X}+C_{X}^{\prime}R_{f}^{2},\gamma_{2}=C_{X}+C_{X}^{\prime}R_{f},\gamma_{3}=\frac{C_{s}R_{f}^{3}}{T_{0}^{3}}$.
@@
$\bar{g}_{i}(x^{\prime},w^{\prime},t^{\prime})=\sum_{u,j}g_{i}(u/N_{1},j/N_{2})\Psi_{u,j}(x^{\prime},w^{\prime},t^{\prime}),$ (B.10)
@@
$\Psi_{u,j}(x^{\prime},w^{\prime},t^{\prime}):=\psi\big{(}3N_{2}(t^{\prime}-j/N_{2})\big{)}\prod_{r=1}^{d_{x}}\psi\big{(}3N_{1}(x^{\prime}_{r}-u_{r}/N_{1})\big{)}\prod_{r=1}^{d_{y}}\psi\big{(}3N_{1}(w^{\prime}_{r}-u_{r+d_{x}}/N_{1})\big{)},$ (B.11)
$\psi(a):=\left\{\begin{array}[]{ll}1,&|a|<1\\ 2-|a|,&1\leq|a|<2\\ 0,&|a|>\geq 2\end{array}\right.$ (B.12)
@@
$\sup_{o^{\prime},t^{\prime}}\Big{|}\bar{g}_{i}(o^{\prime},t^{\prime})-g_{i}(o^{\prime},t^{\prime})|$ $\leq\sup_{o^{\prime},t^{\prime}}\Big{|}\sum_{u,j}(g_{i}(\frac{u}{N_{1}},\frac{j}{N_{2}})-g_{i}(o^{\prime},t^{\prime}))\Psi_{u,j}(o^{\prime},t^{\prime})\Big{|}$ (B.13)
@@
$\phi_{\text{square}}^{l}(a)=a-\sum_{m=1}^{l}2^{-2m}r_{m}(a),\ r_{m}=\underbrace{r\circ\cdots\circ r}_{m\text{ compositions}}$ (B.14)
$\phi_{\text{mul}}^{l}(a,b)=\phi_{\text{square}}^{l}(\frac{a+b}{2})-\phi_{\text{square}}^{l}(\frac{a-b}{2})$ (B.15)
@@
$\widehat{\Psi}_{u,j}(x^{\prime},w^{\prime},t^{\prime}):=\phi_{\text{mul}}^{l}\left(\psi\big{(}3N_{2}(t^{\prime}-j/N_{2})\big{)},\phi_{\text{mul}}^{l}\left(\psi\big{(}3N_{1}(x^{\prime}_{1}-u_{1}/N_{2})\big{)},\cdots\right)\right)$ (B.17)
@@
$\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime}):=\sum_{u,j}g_{i}(u/N_{1},j/N_{2})\widehat{\Psi}_{u,j}(x^{\prime},w^{\prime},t^{\prime}).$ (B.18)
@@
$\|\widehat{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}$ $\leq\|\widehat{g}_{i}-\bar{g}_{i}\|_{L^{\infty}(\Omega)}+|\bar{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}$ (B.19)
@@
$|\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime})-\widehat{g}_{i}(x^{\prime},w^{\prime\prime},t^{\prime})|\lesssim(d_{x}+d_{y})\left(\gamma_{w}+N_{1}\|g_{i}\|_{L^{\infty}(\Omega)}2^{-l+d_{x}+d_{y}}\right)\|w^{\prime}-w^{\prime\prime}\|_{\infty}.$ (B.20)
@@
$\|\widehat{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}\leq\varepsilon/2,\ |\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime})-\widehat{g}_{i}(x^{\prime},w^{\prime\prime},t^{\prime})|\lesssim\gamma_{w}(d_{x}+d_{y})\|w^{\prime}-w^{\prime\prime}\|_{\infty}.$ (B.21)
@@
$\|\widehat{f}-f\|_{L^{\infty}(\Omega_{R_{f}})}\leq\|\widehat{g}-g\|_{L^{\infty}(\Omega)}\leq\sqrt{d_{x}}\varepsilon/2,\ \text{and }\widehat{f}(x,w,t)=0,\forall\ \|x\|_{\infty}>R_{f}.$ (B.22)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h_{*}(y),t)-f_{*}^{\mathbb{P}}(x_{t},h_{*}(y),t)\|^{2}]$ $\leq\mathbb{E}_{t,x_{t},y}\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}\cdot\mathbbm{1}_{\|x_{t}\|\leq R_{f}}$ (B.23)

### v2

@@ B.4 Auxiliary Lemmas
$\displaystyle\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big|\lesssim\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK}}+\frac{b\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK},$ (B.120)
$\displaystyle\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\varphi(x_{i}^{k}))^{2}-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi^{2}]\Big|\lesssim b\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK}}+\frac{b^{2}\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK}.$ (B.121)
@@
$\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big|\lesssim\mathcal{R}_{K,n}(\Phi^{(j_{0}:)})+\sqrt{\frac{(B\epsilon_{j_{0}}+B_{0})\log\big((\log nK)/\delta\big)}{nK}}+\frac{b\log\big((\log nK)/\delta\big)}{nK}.$ (B.122)
From now on we reason on the conjunction of (B.120), (B.121) and (B.122). Define
$U_{j}=B\epsilon_{j}+B_{0}+b\mathcal{R}_{K,n}(\Phi^{(j)})+\sqrt{\frac{b^{2}(B\epsilon_{j}+B_{0})\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK}}+\frac{b^{2}\log\big(\log(b/\epsilon_{j})/\delta\big)}{nK}.$ (B.123)
and thus for any $\varphi\in\Phi^{(j)}$, we have $\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}(\varphi(x_{i}^{k}))^{2}\leq CU_{j}$ for some absolute constant $C$ by (B.121), indicating that $\mathcal{R}_{K,n}(\Phi^{(j)})\leq\phi_{K,n}(CU_{j})\leq\sqrt{C}\phi_{K,n}(U_{j})$. For any $j\leq j_{0}$,
$U_{j}\leq 2(B\epsilon_{j}+B_{0})+b\sqrt{C}\phi_{K,n}(U_{j})+\frac{2b^{2}\log\big((\log nK)/\delta\big)}{nK}.$ (B.124)
@@
$U_{j}\lesssim b^{2}r_{K,n}^{*}+B\epsilon_{j}+B_{0}+\frac{b^{2}\log\big((\log nK)/\delta\big)}{nK}=:r_{K,n}(\epsilon_{j}).$ (B.125)
Therefore, for any $\varphi\in\Phi^{(j)},j\leq j_{0}$, by (B.120),
$\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big|$ $\lesssim\phi_{K,n}(r_{K,n}(\epsilon_{j}))+\sqrt{\frac{(B\epsilon_{j}+B_{0})\log\big((\log nK)/\delta\big)}{nK}}+\frac{b\log\big((\log nK)/\delta\big)}{nK}$ (B.126)
@@
$\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big|\lesssim F_{K,n}(\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]).$ (B.127)
@@
$\displaystyle\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{2}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big((\log nK)/\delta\big)}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log nK)/\delta\big)}{nK}},$ (B.128)
$\displaystyle\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})\leq 2\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]+C^{\prime}(B\vee b)\left(r_{K,n}^{*}+\frac{\log\big((\log nK)/\delta\big)}{nK}\right)+C^{\prime}\sqrt{\frac{B_{0}\log\big((\log nK)/\delta\big)}{nK}}.$ (B.129)
Here $C^{\prime}$ is an absolute constant. Moreover, when $\varphi\in\Phi^{(j)}$ for $j>j_{0}$, we have $\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\leq\frac{b}{nK}$, and according to (B.122),
$\Big|\frac{1}{nK}\sum_{k=1}^{K}\sum_{i=1}^{n}\varphi(x_{i}^{k})-\mathbb{E}_{\widehat{\mathbb{P}}^{(K)}}[\varphi]\Big|\lesssim F_{K,n}(\varepsilon_{j_{0}}).$ (B.130)
@@
B.5 Verifying Task Diversity Assumption
@@
Proposition B.13.
[…]
$\widetilde{\Delta}=2\mathbb{E}_{(x,y)\sim\mathbb{P}^{0}}\mathbb{E}_{t,x_{t}}\left[\left\|\frac{1}{\lambda}\sum_{k=1}^{K}\lambda_{k}\nabla\log p_{t}^{k}(x_{t}|y)-\nabla\log p_{t}^{0}(x_{t}|y)\right\|^{2}\right].$ (B.131)
@@
$L^{\mathbb{P}^{0}}(s_{\widetilde{f}^{0},h})$ $=\mathbb{E}_{\mathbb{P}^{0}}\left[\|\widetilde{f}^{0}(x_{t},h(y),t)-\nabla\log p_{t}^{0}(x_{t}|y)\|^{2}\right]$ (B.132)
@@
Appendix C Proofs in Section 4
C.1 Proofs of Score Network Approximation
Theorem C.1 (Thm. 4.1).
@@ C.1 Proofs of Score Network Approximation
$\displaystyle\inf_{h\in\mathcal{H}}\frac{1}{K}\sum_{k=1}^{K}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}^{k}}[\ell^{\mathbb{P}^{k}}(x,y,s_{f,h})]=\mathcal{O}\left(\log^{2}(nK/(\varepsilon\delta))\varepsilon^{2}\right),\text{ (transfer learning)}$ (C.1)
$\displaystyle\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]=\mathcal{O}\left(\log^{2}(nK/(\varepsilon\delta))\varepsilon^{2}\right),\text{ (meta-learning)}$ (C.2)
@@
$L_{f}=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),W_{f}=\mathcal{O}\left(\frac{\log^{3(d_{x}+d_{y})/2}(nK/(\varepsilon\delta))}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right),$ (C.3)
@@
$L_{h}=\mathcal{O}\left(\log(1/\varepsilon)\right),W_{h}=\mathcal{O}\left(\varepsilon^{-D_{y}}\log(1/\varepsilon)\right),$ (C.4)
@@
$\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ (C.5)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}]$ $\leq\mathbb{E}_{t,x_{t},y}\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}\cdot\mathbbm{1}_{\|x_{t}\|\leq R_{f}}$ (C.6)
@@
where $\Omega_{R_{f}}=[-R_{f},R_{f}]^{d_{x}}\times[0,1]^{d_{y}}\times[T_{0},T]$. By Proposition C.2, C.3,
$\inf_{h\in\mathcal{H}}\mathbb{E}_{\mathbb{P}\sim\mathbb{P}_{\text{meta}}}\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}[\ell^{\mathbb{P}}(x,y,s_{f,h})]$ (C.7)
[…]
@@
$\inf_{f\in\mathcal{F}}\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h_{*}(y),t)-f^{\mathbb{P}}(x_{t},h_{*}(y),t)\|^{2}]\leq d_{x}\varepsilon^{2},$ (C.8)
@@
$L_{f}=\mathcal{O}\left(\log\left(\frac{\log(nK/(\varepsilon\delta))}{\varepsilon}\right)\right),W_{f}=\mathcal{O}\left(\frac{\log^{3(d_{x}+d_{y})/2}(nK/(\varepsilon\delta))}{\varepsilon^{d_{x}+d_{y}+1}T_{0}^{3}}\right),$ (C.9)
@@
By Lemma B.2 and C.6, $f_{*}$ is $\gamma_{1}$-Lipschitz in $x$, $\gamma_{2}$-Lipschitz in $w$, and $\gamma_{3}$-Lipshcitz in $t$ in a bounded domain $\Omega_{R_{f}}$, where $\gamma_{1}=C_{X}+C_{X}^{\prime}R_{f}^{2},\gamma_{2}=C_{X}+C_{X}^{\prime}R_{f},\gamma_{3}=\frac{C_{s}R_{f}^{3}}{T_{0}^{3}}$.
@@
$\bar{g}_{i}(x^{\prime},w^{\prime},t^{\prime})=\sum_{u,j}g_{i}(u/N_{1},j/N_{2})\Psi_{u,j}(x^{\prime},w^{\prime},t^{\prime}),$ (C.10)
@@
$\Psi_{u,j}(x^{\prime},w^{\prime},t^{\prime}):=\psi\big(3N_{2}(t^{\prime}-j/N_{2})\big)\prod_{r=1}^{d_{x}}\psi\big(3N_{1}(x^{\prime}_{r}-u_{r}/N_{1})\big)\prod_{r=1}^{d_{y}}\psi\big(3N_{1}(w^{\prime}_{r}-u_{r+d_{x}}/N_{1})\big),$ (C.11)
$\psi(a):=\left\{\begin{array}[]{ll}1,&|a|<1\\ 2-|a|,&1\leq|a|<2\\ 0,&|a|>\geq 2\end{array}\right$. (C.12)
@@
$\sup_{o^{\prime},t^{\prime}}\Big|\bar{g}_{i}(o^{\prime},t^{\prime})-g_{i}(o^{\prime},t^{\prime})|$ $\leq\sup_{o^{\prime},t^{\prime}}\Big|\sum_{u,j}(g_{i}(\frac{u}{N_{1}},\frac{j}{N_{2}})-g_{i}(o^{\prime},t^{\prime}))\Psi_{u,j}(o^{\prime},t^{\prime})\Big|$ (C.13)
@@
$\phi_{\text{square}}^{l}(a)=a-\sum_{m=1}^{l}2^{-2m}r_{m}(a),\ r_{m}=\underbrace{r\circ\cdots\circ r}_{m\text{ compositions}}$ (C.14)
$\phi_{\text{mul}}^{l}(a,b)=\phi_{\text{square}}^{l}(\frac{a+b}{2})-\phi_{\text{square}}^{l}(\frac{a-b}{2})$ (C.15)
@@
$\widehat{\Psi}_{u,j}(x^{\prime},w^{\prime},t^{\prime}):=\phi_{\text{mul}}^{l}\left(\psi\big(3N_{2}(t^{\prime}-j/N_{2})\big),\phi_{\text{mul}}^{l}\left(\psi\big(3N_{1}(x^{\prime}_{1}-u_{1}/N_{2})\big),\cdots\right)\right)$ (C.17)
@@
$\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime}):=\sum_{u,j}g_{i}(u/N_{1},j/N_{2})\widehat{\Psi}_{u,j}(x^{\prime},w^{\prime},t^{\prime}).$ (C.18)
@@
$\|\widehat{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}$ $\leq\|\widehat{g}_{i}-\bar{g}_{i}\|_{L^{\infty}(\Omega)}+|\bar{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}$ (C.19)
@@
$|\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime})-\widehat{g}_{i}(x^{\prime},w^{\prime\prime},t^{\prime})|\lesssim(d_{x}+d_{y})\left(\gamma_{w}+N_{1}\|g_{i}\|_{L^{\infty}(\Omega)}2^{-l+d_{x}+d_{y}}\right)\|w^{\prime}-w^{\prime\prime}\|_{\infty}.$ (C.20)
@@
$\|\widehat{g}_{i}-g_{i}\|_{L^{\infty}(\Omega)}\leq\varepsilon/2,\ |\widehat{g}_{i}(x^{\prime},w^{\prime},t^{\prime})-\widehat{g}_{i}(x^{\prime},w^{\prime\prime},t^{\prime})|\lesssim\gamma_{w}(d_{x}+d_{y})\|w^{\prime}-w^{\prime\prime}\|_{\infty}.$ (C.21)
@@
$\|\widehat{f}-f\|_{L^{\infty}(\Omega_{R_{f}})}\leq\|\widehat{g}-g\|_{L^{\infty}(\Omega)}\leq\sqrt{d_{x}}\varepsilon/2,\ \text{and }\widehat{f}(x,w,t)=0,\forall\ \|x\|_{\infty}>R_{f}.$ (C.22)
@@
$\mathbb{E}_{(x,y)\sim\mathbb{P}}\mathbb{E}_{t,x_{t}|x}[\|f(x_{t},h_{*}(y),t)-f_{*}^{\mathbb{P}}(x_{t},h_{*}(y),t)\|^{2}]$ $\leq\mathbb{E}_{t,x_{t},y}\|f(x_{t},h(y),t)-f(x_{t},h_{*}(y),t)\|^{2}\cdot\mathbbm{1}_{\|x_{t}\|\leq R_{f}}$ (C.23)